OpenStudy (henrietepurina):

A machine used to make badges costs $150. It costs $2 to make each badge. Part A: Write a function to show the total cost f(n) of making n badges. (4 points) Part B: What is the total cost of making 20 badges? (4 points) Part C: If the cost of the machine was $20 more, what would be the cost of making 7 badges? (2 points)

OpenStudy (yanasidlinskiy):

It would be the initial cost of the machine, plus the cost to make each badge so:150+2n. where n is the number of badges.

OpenStudy (henrietepurina):

So the answer for A is?

OpenStudy (yanasidlinskiy):

And then from there for Part B it is 120 + 2(20) since it is asked for 20 badge

OpenStudy (paki):

hm... see... Part A will be... let x = number of badges C(x) = cost of making x number of badges function: C(x) = 2x + 150 Part B will be... C(20) = 2(20) + 150 = $190 Part C will be.. 2(7) + 170 = $184 @HenrietePurina

OpenStudy (yanasidlinskiy):

YOU have no choices. You have Part A, Part B, and Part C.

OpenStudy (henrietepurina):

thanks alot guys!(or girls)

OpenStudy (paki):

i am boy :-)

OpenStudy (henrietepurina):

i was talking about yanasidlinskiy

OpenStudy (yanasidlinskiy):

Oh. thanx:) Your Welcome!:) No problem!:)

OpenStudy (paki):

ahhh cool... ok :-) @HenrietePurina

OpenStudy (henrietepurina):

i love people that help immediatly so because of this i have you 2 will have a new fan

OpenStudy (henrietepurina):

why did this comment go here? ^

OpenStudy (paki):

ahhh thanks HenrietePurina

OpenStudy (henrietepurina):

you too

OpenStudy (paki):

pleasure